# Latent Factor Analysis in Short Panels

### v2

@@ B Proofs of Propositions 1-5
Let us now establish the asymptotic normality of $vech(Z^{*}_{n})$. For any integer $m$, we let $A_{m}$ denote the unique $m^{2}\times\frac{m(m+1)}{2}$ matrix satisfying $vec(S)=A_{m}vech(S)$ for any $m\times m$ symmetric matrix $S$ [^34] The explicit form for $A_{m}$ is $A_{m}=\left[\sqrt{2}(e_{1}\otimes e_{1}):\cdots:\sqrt{2}(e_{m}\otimes e_{m}):\{e_{i}\otimes e_{j}+e_{j}\otimes e_{i}\}_{i<j}\right]$, with $e_{i}$ being the $i$th unit vector of dimension $m$. Matrix $A_{m}$ satisfies $A_{m}^{\prime}A_{m}=2I_{\frac{m(m+1)}{2}}$, $A_{m}A_{m}^{\prime}=I_{m^{2}}+K_{m,m}$, and $K_{m,m}A_{m}=A_{m}$, where $K_{m,m}$ is the commutation matrix (see also Magnus, Neudecker (2007) Theorem 12 in Chapter 2.8). Then, we have $vech(Z^{*}_{n})=\boldsymbol{R}^{\prime}vech(\mathscr{Z}_{n})$, where $\mathscr{Z}_{n}=V_{\varepsilon}^{-1/2}Z_{n}V_{\varepsilon}^{-1/2}$, $\boldsymbol{R}=\frac{1}{2}A^{\prime}_{T}(Q\otimes Q)A_{T-k}$, and $Q=V_{\varepsilon}^{-1/2}G$. Matrix $\boldsymbol{R}$ satisfies $\boldsymbol{R}^{\prime}\boldsymbol{R}=I_{p}$. The next lemma establishes the asymptotic normality of $vech(\mathscr{Z}_{n})$.
@@
(a) Under Assumptions 1-2, A.2, A.6 (a)-(b), we have $\Omega_{n}^{-1/2}vech(\mathscr{Z}_{n})\Rightarrow N(0,I_{\frac{T(T+1)}{2}})$ as $n\rightarrow\infty$ and $T$ is fixed, where $\Omega_{n}=D_{n}+\kappa_{n}I_{\frac{T(T+1)}{2}}$, and $\kappa_{n}=\frac{1}{n}\sum_{m=1}^{J_{n}}\left(\sum_{i\neq j\in I_{m}}\sigma_{ij}^{2}\right)$. If additionally Assumption A.6 (c) holds, then $vech(\mathscr{Z}_{n})\Rightarrow N(0,\Omega)$, with $\Omega:=D+\kappa I_{\frac{T(T+1)}{2}}$.
@@
(b) We have $\hat{z}^{*}_{m,n}=\sum_{i\in I_{m}}\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}\left(\tilde{y}_{i}\tilde{y}_{i}^{\prime}\right)\hat{V}_{\varepsilon}^{-1}\hat{G}$ with $\tilde{y}_{i}=y_{i}-\bar{y}$, since $\hat{\varepsilon}_{i}=M_{\hat{F},\hat{V}_{\varepsilon}}\tilde{y}_{i}$ and $\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}M_{\hat{F},\hat{V}_{\varepsilon}}=\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}$. We get $\hat{z}^{*}_{m,n}=\sum_{i\in I_{m}}\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}\left(\tilde{\varepsilon}_{i}\tilde{\varepsilon}_{i}^{\prime}\right)\hat{V}_{\varepsilon}^{-1}\hat{G}+\sum_{i\in I_{m}}\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}\left(F\beta_{i}\beta_{i}^{\prime}F^{\prime}\right)\hat{V}_{\varepsilon}^{-1}\hat{G}\break+\sum_{i\in I_{m}}\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}\left(F\beta_{i}\tilde{\varepsilon}_{i}^{\prime}+\tilde{\varepsilon}_{i}\beta_{i}^{\prime}F^{\prime}\right)\hat{V}_{\varepsilon}^{-1}\hat{G}=:\tilde{z}^{*}_{m,n}+z^{*}_{m,n,1}+z^{*}_{m,n,2}$, where $\tilde{\varepsilon}_{i}=\varepsilon_{i}-\bar{\varepsilon}$ by using $\tilde{y}_{i}\tilde{y}_{i}^{\prime}=\tilde{\varepsilon}_{i}\tilde{\varepsilon}_{i}^{\prime}+F\beta_{i}\beta_{i}^{\prime}F^{\prime}+F\beta_{i}\tilde{\varepsilon}_{i}^{\prime}+\tilde{\varepsilon}_{i}\beta_{i}^{\prime}F^{\prime}$. Then, we can decompose $\hat{\Omega}_{Z^{*}}$ into a sum of a leading term and other terms, which are asymptotically negligible, so that $\hat{\Omega}_{Z^{*}}=\tilde{\Omega}_{Z^{*}}+o_{p}(1)$, with $\tilde{\Omega}_{Z^{*}}=\frac{1}{n}\sum_{m=1}^{J_{n}}vech(\bar{z}^{*}_{m,n})vech(\bar{z}^{*}_{m,n})^{\prime}$, with $\bar{z}^{*}_{m,n}$ defined as $\tilde{z}^{*}_{m,n}$ after replacing $\tilde{\varepsilon}_{i}$ with $\varepsilon_{i}$. Let us now show that $M_{\hat{\boldsymbol{X}}}\tilde{\Omega}_{Z^{*}}M_{\hat{\boldsymbol{X}}}=M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}+o_{p}(1)$ up to pre- and post-multiplication by a rotation matrix and its inverse. We have $M_{\hat{\boldsymbol{X}}}vech\left(\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}\left(\varepsilon_{i}\varepsilon_{i}^{\prime}\right)\hat{V}_{\varepsilon}^{-1}\hat{G}\right)=M_{\hat{\boldsymbol{X}}}vech\left(\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}\left(\varepsilon_{i}\varepsilon_{i}^{\prime}-\sigma_{ii}V_{\varepsilon}\right)\hat{V}_{\varepsilon}^{-1}\hat{G}\right)$, because of the kernel of $M_{\hat{\boldsymbol{X}}}$ Moreover, from the properties of matrix $A_{m}$ introduced in part (a), we have: $vech\left(\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}(\varepsilon_{i}\varepsilon_{i}^{\prime}-\sigma_{ii}V_{\varepsilon})\hat{V}_{\varepsilon}^{-1}\hat{G}\right)\break=vech\left(\tilde{Q}^{\prime}(e_{i}e_{i}^{\prime}-\sigma_{ii}I_{T})\tilde{Q}\right)=\hat{\boldsymbol{R}}^{\prime}vech(e_{i}e_{i}^{\prime}-\sigma_{ii}I_{T}),$ where $e_{i}=V_{\varepsilon}^{-1/2}\varepsilon_{i}$, and $\hat{\boldsymbol{R}}:=\frac{1}{2}A_{T}^{\prime}(\tilde{Q}\otimes\tilde{Q})A_{T-k}$ with $\tilde{Q}=V_{\varepsilon}^{1/2}\hat{V}_{\varepsilon}^{-1}\hat{G}$. We get $M_{\hat{\boldsymbol{X}}}vech(\bar{z}^{*}_{m,n})={M}_{\hat{\boldsymbol{X}}}\hat{\boldsymbol{R}}^{\prime}vech(\zeta_{m,n})$, where $\zeta_{m,n}:=\sum_{i\in I_{m}}(e_{i}e_{i}^{\prime}-\sigma_{ii}I_{T})$. Besides, $vech(\mathscr{Z}_{n})=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}vech(\zeta_{m,n})$. Then, $M_{\hat{\boldsymbol{X}}}\tilde{\Omega}_{Z^{*}}M_{\hat{\boldsymbol{X}}}\break={M}_{\hat{\boldsymbol{X}}}\hat{\boldsymbol{R}}^{\prime}\tilde{\Omega}_{n}\hat{\boldsymbol{R}}{M}_{\hat{\boldsymbol{X}}}$ for $\tilde{\Omega}_{n}:=\frac{1}{n}\sum_{m=1}^{J_{n}}vech(\zeta_{m,n})vech(\zeta_{m,n})^{\prime}$. Further, $E[\tilde{\Omega}_{n}]=V[vech(\mathscr{Z}_{n})]\break=\Omega_{n}$. Moreover, $\tilde{\Omega}_{n}-E[\tilde{\Omega}_{n}]=o_{p}(1)$, by using $vec(\tilde{\Omega}_{n})=\frac{1}{n}\sum_{m=1}^{J_{n}}vech(\zeta_{m,n})\otimes vech(\zeta_{m,n})$ and $\|V[vec(\tilde{\Omega}_{n})]\|\leq C\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E\left[\|vech(\zeta_{m,n})\|^{4}\right]=o(1)$, where the latter bound is shown in the proof of Lemma 1 using Assumption 2 (d). Additionally, by Assumption A.6, we have $\Omega_{n}=\Omega+o(1)$. Thus, $\tilde{\Omega}_{n}=\Omega+o_{p}(1)$. Now, from the proof of part (a) we have $\hat{G}\hat{O}=G+o_{p}(1)$ for some $(T-k)\times(T-k)$ orthogonal matrix $\hat{O}$. Then, by Proposition 9 (e) in Appendix E.6, we have $\hat{\boldsymbol{R}}{M}_{\hat{\boldsymbol{X}}}\hat{\mathscr{R}}^{-1}=\boldsymbol{R}M_{\boldsymbol{X}}+o_{p}(1)$, for a $p$ dimensional orthogonal matrix ^$\hat{\mathscr{R}}\equiv\mathscr{R}(\hat{O})$. We conclude that ^$\hat{\mathscr{R}}M_{\hat{\boldsymbol{X}}}\tilde{\Omega}_{Z^{*}}M_{\hat{\boldsymbol{X}}}\hat{\mathscr{R}}^{-1}$ is a consistent estimator of $M_{\boldsymbol{X}}{\Omega}_{Z^{*}}M_{\boldsymbol{X}}$ as $n\rightarrow\infty$ and $T$ is fixed. Part (b) then follows from the continuity of eigenvalues for symmetric matrices, and their invariance under pre- and post-multiplication by an orthogonal matrix and its transpose.
@@
Proof of Proposition 4: Let us first get the asymptotic expansion of $\hat{V}_{y}-\tilde{V}_{\varepsilon}=\frac{1}{n}\tilde{Y}\tilde{Y}^{\prime}-\tilde{V}_{\varepsilon}$. With the drifting DGP $Y=\mu 1_{n}^{\prime}+F\beta^{\prime}+F_{k+1}\beta_{loc}^{\prime}+\varepsilon$, and using $\bar{\beta}=0$, $\bar{\beta}_{loc}=0$, $\frac{1}{n}[\beta:\beta_{loc}]^{\prime}[\beta:\beta_{loc}]=I_{k+1}$ and Lemma 5 (a) in Appendix D, we get $\hat{V}_{y}-\tilde{V}_{\varepsilon}=FF^{\prime}+\frac{1}{\sqrt{n}}\Psi_{y,loc}+R_{y}$, where
$\Psi_{y,loc}=c_{k+1}\rho_{k+1}\rho_{k+1}^{\prime}+\frac{1}{\sqrt{n}}(\varepsilon\beta F^{\prime}+F\beta^{\prime}\varepsilon^{\prime})+\sqrt{n}\left(\frac{1}{n}\varepsilon\varepsilon^{\prime}-\tilde{V}_{\varepsilon}\right),$ (B.3)
@@
$f(z;\lambda_{1},...,\lambda_{df})=\sum_{k=0}^{\infty}\bar{c}_{k}(\lambda_{1},...,\lambda_{df})g(z;df+2k,0),$ (B.4)
@@
where the coefficients $\bar{c}_{k}(\lambda_{1},...,\lambda_{df})=Ae^{-\sum_{j=1}^{df}\lambda_{j}^{2}/2}E[Q(\lambda_{1},...,\lambda_{df})^{k}]/k!$ involve moments of the quadratic form $Q(\lambda_{1},...,\lambda_{df})=(1/2)\sum_{j=1}^{df}\left(\nu_{j}^{1/2}X_{j}+\lambda_{j}(1-\nu_{j})^{1/2}\right)^{2}$ of the mutually independent variables $X_{j}\sim N(0,1)$, $A=\prod_{j=1}^{df}\mu_{j}^{-1/2}$, and $\nu_{j}=1-\frac{1}{\mu_{j}}\min_{\ell}\mu_{\ell}$. Without loss of generality for checking the monotonicity, we have rescaled the density so that $\min_{j}\mu_{j}=1$. Then, from (B.4), we get the ratio: $\frac{f(z;\lambda_{1},...,\lambda_{df})}{f(z;0,...,0)}=\frac{\sum_{k=0}^{\infty}\bar{c}_{k}(\lambda_{1},...,\lambda_{df})g(z;df+2k,0)}{\sum_{k=0}^{\infty}\bar{c}_{k}(0,...,0)g(z;df+2k,0)}.$ By dividing both the numerator and the denominator by the central chi-square density $g(z;df,0)$, we get $\frac{f(z;\lambda_{1},...,\lambda_{df})}{f(z;0,...,0)}=e^{-\sum_{j=1}^{df}\lambda_{j}^{2}/2}\frac{\sum_{k=0}^{\infty}c_{k}(\lambda_{1},...,\lambda_{df})\psi_{k}(z)}{\sum_{k=0}^{\infty}c_{k}(0,...,0)\psi_{k}(z)}=:e^{-\sum_{j=1}^{df}\lambda_{j}^{2}/2}\Psi(z;\lambda_{1},...,\lambda_{df}),$ where $\psi_{k}(z):=$ $g(z;df+2k,0)/g(z;df,0)=\frac{\Gamma(\frac{df}{2})}{2^{k}\Gamma(\frac{df}{2}+k)}z^{k}$ is the ratio of central chi-square distributions with $df+2k$ and $df$ degrees of freedom, and $c_{k}(\lambda_{1},...,\lambda_{df})=E[Q(\lambda_{1},...,\lambda_{df})^{k}]/k!$. We use the short notation $c_{k}(\lambda):=c_{k}(\lambda_{1},...,\lambda_{df})$ and $c_{k}(0):=c_{k}(0,...,0)$. The factor $e^{-\sum_{j=1}^{df}\lambda_{j}^{2}/2}$ does not impact on the monotonicity of the density ratio. We take the derivative of $\Psi(z;\lambda_{1},...,\lambda_{df})$ with respect to argument $z$ and get $\partial_{z}\Psi(z;\lambda_{1},...,\lambda_{df})=\frac{\left(\sum_{k=1}^{\infty}c_{k}(\lambda)\psi_{k}^{\prime}(z)\right)\left(1+\sum_{k=1}^{\infty}c_{k}(0)\psi_{k}(z)\right)}{\left(\sum_{k=0}^{\infty}c_{k}(0)\psi_{k}(z)\right)^{2}}-\frac{\left(1+\sum_{k=1}^{\infty}c_{k}(\lambda)\psi_{k}(z)\right)\left(\sum_{k=1}^{\infty}c_{k}(0)\psi_{k}^{\prime}(z)\right)}{\left(\sum_{k=0}^{\infty}c_{k}(0)\psi_{k}(z)\right)^{2}}.$ The sign is given by the difference of the numerators, which is $\sum_{k=1}^{\infty}[c_{k}(\lambda)-c_{k}(0)]\psi_{k}^{\prime}(z)+\sum_{k,l=1,k\neq l}^{\infty}c_{k}(\lambda)c_{l}(0)[\psi_{k}^{\prime}(z)\psi_{l}(z)-\psi_{k}(z)\psi_{l}^{\prime}(z)]=\sum_{k=1}^{\infty}[c_{k}(\lambda)-c_{k}(0)]\psi_{k}^{\prime}(z)+\sum_{k,l=1,k>l}^{\infty}[c_{k}(\lambda)c_{l}(0)-c_{l}(\lambda)c_{k}(0)][\psi_{k}^{\prime}(z)\psi_{l}(z)-\psi_{k}(z)\psi_{l}^{\prime}(z)].$ We use $\psi_{k}^{\prime}(z)=\frac{\Gamma(\frac{d}{2})k}{2^{k}\Gamma(\frac{d}{2}+k)}z^{k-1}$ and $\psi_{k}^{\prime}(z)\psi_{l}(z)-\psi_{k}(z)\psi_{l}^{\prime}(z)=(k-l)\frac{\Gamma(\frac{d}{2})^{2}}{2^{k+l}\Gamma(\frac{d}{2}+k)\Gamma(\frac{d}{2}+l)}z^{k+l-1}$ for $k>l$ and $z\geq 0$. The difference of the numerators in the derivative of the density ratio becomes: $\frac{1}{2}\frac{\Gamma(\frac{d}{2})}{\Gamma(\frac{d}{2}+1)}[c_{1}(\lambda)-c_{1}(0)]+\frac{1}{2^{2}}\frac{2\Gamma(\frac{d}{2})}{\Gamma(\frac{d}{2}+2)}[c_{2}(\lambda)-c_{2}(0)]z+\sum_{m=3}^{\infty}\frac{1}{2^{m}}\left(m\frac{\Gamma(\frac{d}{2})}{\Gamma(\frac{d}{2}+m)}[c_{m}(\lambda)-c_{m}(0)]\right.$ $\left.+\sum_{k>l\geq 1,k+l=m}\frac{(k-l)\Gamma(\frac{d}{2})^{2}}{\Gamma(\frac{d}{2}+k)\Gamma(\frac{d}{2}+l)}[c_{k}(\lambda)c_{l}(0)-c_{l}(\lambda)c_{k}(0)]\right)z^{m-1}=\sum_{m=1}^{\infty}\frac{1}{2^{m}}\kappa_{m}z^{m-1},$ with $\kappa_{m}:=\sum_{k>l\geq 0,k+l=m}(k-l)\frac{\Gamma(\frac{d}{2})^{2}}{\Gamma(\frac{d}{2}+k)\Gamma(\frac{d}{2}+l)}$ $[c_{k}(\lambda)c_{l}(0)-c_{l}(\lambda)c_{k}(0)]$. A direct calculation shows that $\kappa_{1},\kappa_{2}\geq 0$. Hence, a sufficient condition for monotonicity of the density ratio is $\kappa_{m}\geq 0$, for all $m\geq 3$, i.e., Inequalities (7). Thus, the test rejects for large values of the argument, i.e., $\phi(z)=\boldsymbol{1}\{z\geq\bar{C}\}$, where the constant $\bar{C}$ is determined by fixing the asymptotic size under the null hypothesis.

### v3

@@ A Proofs
(iii) Let us now establish the asymptotic normality of $vech(Z^{*}_{n})$. For any integer $m$, we let $A_{m}$ denote the unique $m^{2}\times\frac{m(m+1)}{2}$ matrix satisfying $vec(S)=A_{m}vech(S)$ for any $m\times m$ symmetric matrix SS. [^7] The explicit form for $A_{m}$ is $A_{m}=\left[\sqrt{2}(e_{1}\otimes e_{1}):\cdots:\sqrt{2}(e_{m}\otimes e_{m}):\{e_{i}\otimes e_{j}+e_{j}\otimes e_{i}\}_{i<j}\right]$, with $e_{i}$ being the $i$th unit vector of dimension $m$. Duplication matrix $A_{m}$ satisfies $A_{m}^{\prime}A_{m}=2I_{\frac{m(m+1)}{2}}$, $A_{m}A_{m}^{\prime}=I_{m^{2}}+K_{m,m}$, and $K_{m,m}A_{m}=A_{m}$, where $K_{m,m}$ is the commutation matrix (see also Magnus, Neudecker (2007) Theorem 12 in Chapter 2.8). Then, we have $vech(Z^{*}_{n})=\boldsymbol{R}^{\prime}vech(\mathscr{Z}_{n})$, where $\mathscr{Z}_{n}=V_{\varepsilon}^{-1/2}Z_{n}V_{\varepsilon}^{-1/2}$, $\boldsymbol{R}=\frac{1}{2}A^{\prime}_{T}(Q\otimes Q)A_{T-k}$, and $Q=V_{\varepsilon}^{-1/2}G$. Matrix $\boldsymbol{R}$ satisfies $\boldsymbol{R}^{\prime}\boldsymbol{R}=I_{p}$. The next lemma establishes the asymptotic normality of $vech(\mathscr{Z}_{n})$.
@@
(a) Under Assumptions 1-2, 4, 8 (a)-(b), we have $\Omega_{n}^{-1/2}vech(\mathscr{Z}_{n})\Rightarrow N(0,I_{\frac{T(T+1)}{2}})$ as $n\rightarrow\infty$ and $T$ is fixed, where $\Omega_{n}=D_{n}+\kappa_{n}I_{\frac{T(T+1)}{2}}$, and $\kappa_{n}=\frac{1}{n}\sum_{m=1}^{J_{n}}\left(\sum_{i\neq j\in I_{m}}\sigma_{ij}^{2}\right)$. If additionally Assumption 8 (c) holds, then $vech(\mathscr{Z}_{n})\Rightarrow N(0,\Omega)$, with $\Omega:=D+\kappa I_{\frac{T(T+1)}{2}}$.
@@
[…]
Proof of Proposition 3: Let us first get the asymptotic expansion of $\hat{V}_{y}-\tilde{V}_{\varepsilon}=\frac{1}{n}\tilde{Y}\tilde{Y}^{\prime}-\tilde{V}_{\varepsilon}$. With the drifting DGP $Y=\mu 1_{n}^{\prime}+F\beta^{\prime}+F_{k+1}\beta_{loc}^{\prime}+\varepsilon$, and using $\bar{\beta}=0$, $\bar{\beta}_{loc}=0$, $\frac{1}{n}[\beta:\beta_{loc}]^{\prime}[\beta:\beta_{loc}]=I_{k+1}$ and Lemma 6 (a) in Appendix B, we get $\hat{V}_{y}-\tilde{V}_{\varepsilon}=FF^{\prime}+\frac{1}{\sqrt{n}}\Psi_{y,loc}+R_{y}$, where
$\Psi_{y,loc}=c_{k+1}\rho_{k+1}\rho_{k+1}^{\prime}+\frac{1}{\sqrt{n}}(\varepsilon\beta F^{\prime}+F\beta^{\prime}\varepsilon^{\prime})+\sqrt{n}\left(\frac{1}{n}\varepsilon\varepsilon^{\prime}-\tilde{V}_{\varepsilon}\right),$ (A.4)
@@
$f(z;\lambda_{1},...,\lambda_{df})=\sum_{k=0}^{\infty}\bar{c}_{k}(\lambda_{1},...,\lambda_{df})g(z;df+2k,0),$ (A.5)
@@
where the coefficients $\bar{c}_{k}(\lambda_{1},...,\lambda_{df})=Ae^{-\sum_{j=1}^{df}\lambda_{j}^{2}/2}E[Q(\lambda_{1},...,\lambda_{df})^{k}]/k!$ involve moments of the quadratic form $Q(\lambda_{1},...,\lambda_{df})=(1/2)\sum_{j=1}^{df}\left(\nu_{j}^{1/2}X_{j}+\lambda_{j}(1-\nu_{j})^{1/2}\right)^{2}$ of the mutually independent variables $X_{j}\sim N(0,1)$, $A=\prod_{j=1}^{df}\mu_{j}^{-1/2}$, and $\nu_{j}=1-\frac{1}{\mu_{j}}\min_{\ell}\mu_{\ell}$. Without loss of generality for checking the monotonicity, we have rescaled the density so that $\min_{j}\mu_{j}=1$. Then, from (A.5), we get the ratio: $\frac{f(z;\lambda_{1},...,\lambda_{df})}{f(z;0,...,0)}=\frac{\sum_{k=0}^{\infty}\bar{c}_{k}(\lambda_{1},...,\lambda_{df})g(z;df+2k,0)}{\sum_{k=0}^{\infty}\bar{c}_{k}(0,...,0)g(z;df+2k,0)}.$ By dividing both the numerator and the denominator by the central chi-square density $g(z;df,0)$, we get $\frac{f(z;\lambda_{1},...,\lambda_{df})}{f(z;0,...,0)}=e^{-\sum_{j=1}^{df}\lambda_{j}^{2}/2}\frac{\sum_{k=0}^{\infty}c_{k}(\lambda_{1},...,\lambda_{df})\psi_{k}(z)}{\sum_{k=0}^{\infty}c_{k}(0,...,0)\psi_{k}(z)}=:e^{-\sum_{j=1}^{df}\lambda_{j}^{2}/2}\Psi(z;\lambda_{1},...,\lambda_{df}),$ where $\psi_{k}(z):=$ $g(z;df+2k,0)/g(z;df,0)=\frac{\Gamma(\frac{df}{2})}{2^{k}\Gamma(\frac{df}{2}+k)}z^{k}$ is the ratio of central chi-square distributions with $df+2k$ and $df$ degrees of freedom, and $c_{k}(\lambda_{1},...,\lambda_{df})=E[Q(\lambda_{1},...,\lambda_{df})^{k}]/k!$. We use the short notation $c_{k}(\lambda):=c_{k}(\lambda_{1},...,\lambda_{df})$ and $c_{k}(0):=c_{k}(0,...,0)$. The factor $e^{-\sum_{j=1}^{df}\lambda_{j}^{2}/2}$ does not impact on the monotonicity of the density ratio. We take the derivative of $\Psi(z;\lambda_{1},...,\lambda_{df})$ with respect to argument $z$ and get $\partial_{z}\Psi(z;\lambda_{1},...,\lambda_{df})=\frac{\left(\sum_{k=1}^{\infty}c_{k}(\lambda)\psi_{k}^{\prime}(z)\right)\left(1+\sum_{k=1}^{\infty}c_{k}(0)\psi_{k}(z)\right)}{\left(\sum_{k=0}^{\infty}c_{k}(0)\psi_{k}(z)\right)^{2}}-\frac{\left(1+\sum_{k=1}^{\infty}c_{k}(\lambda)\psi_{k}(z)\right)\left(\sum_{k=1}^{\infty}c_{k}(0)\psi_{k}^{\prime}(z)\right)}{\left(\sum_{k=0}^{\infty}c_{k}(0)\psi_{k}(z)\right)^{2}}.$ The sign is given by the difference of the numerators, which is $\sum_{k=1}^{\infty}[c_{k}(\lambda)-c_{k}(0)]\psi_{k}^{\prime}(z)+\sum_{k,l=1,k\neq l}^{\infty}c_{k}(\lambda)c_{l}(0)[\psi_{k}^{\prime}(z)\psi_{l}(z)-\psi_{k}(z)\psi_{l}^{\prime}(z)]=\sum_{k=1}^{\infty}[c_{k}(\lambda)-c_{k}(0)]\psi_{k}^{\prime}(z)+\sum_{k,l=1,k>l}^{\infty}[c_{k}(\lambda)c_{l}(0)-c_{l}(\lambda)c_{k}(0)][\psi_{k}^{\prime}(z)\psi_{l}(z)-\psi_{k}(z)\psi_{l}^{\prime}(z)].$ We use $\psi_{k}^{\prime}(z)=\frac{\Gamma(\frac{d}{2})k}{2^{k}\Gamma(\frac{d}{2}+k)}z^{k-1}$ and $\psi_{k}^{\prime}(z)\psi_{l}(z)-\psi_{k}(z)\psi_{l}^{\prime}(z)=(k-l)\frac{\Gamma(\frac{d}{2})^{2}}{2^{k+l}\Gamma(\frac{d}{2}+k)\Gamma(\frac{d}{2}+l)}z^{k+l-1}$ for $k>l$ and $z\geq 0$. The difference of the numerators in the derivative of the density ratio becomes: $\frac{1}{2}\frac{\Gamma(\frac{d}{2})}{\Gamma(\frac{d}{2}+1)}[c_{1}(\lambda)-c_{1}(0)]+\frac{1}{2^{2}}\frac{2\Gamma(\frac{d}{2})}{\Gamma(\frac{d}{2}+2)}[c_{2}(\lambda)-c_{2}(0)]z+\sum_{m=3}^{\infty}\frac{1}{2^{m}}\left(m\frac{\Gamma(\frac{d}{2})}{\Gamma(\frac{d}{2}+m)}[c_{m}(\lambda)-c_{m}(0)]\right$. $\left.+\sum_{k>l\geq 1,k+l=m}\frac{(k-l)\Gamma(\frac{d}{2})^{2}}{\Gamma(\frac{d}{2}+k)\Gamma(\frac{d}{2}+l)}[c_{k}(\lambda)c_{l}(0)-c_{l}(\lambda)c_{k}(0)]\right)z^{m-1}=\sum_{m=1}^{\infty}\frac{1}{2^{m}}\kappa_{m}z^{m-1},$ with $\kappa_{m}:=\sum_{k>l\geq 0,k+l=m}(k-l)\frac{\Gamma(\frac{d}{2})^{2}}{\Gamma(\frac{d}{2}+k)\Gamma(\frac{d}{2}+l)}$ $c_{k}(\lambda)c_{l}(0)-c_{l}(\lambda)c_{k}(0)]$. A direct calculation shows that $\kappa_{1},\kappa_{2}\geq 0$. Hence, a sufficient condition for monotonicity of the density ratio is $\kappa_{m}\geq 0$, for all $m\geq 3$, i.e., Inequalities (5). Thus, the test rejects for large values of the argument, i.e., $\phi(z)=\boldsymbol{1}\{z\geq\bar{C}\}$, where the constant $\bar{C}$ is determined by fixing the asymptotic size under the null hypothesis.